OpenStudy (anonymous):

Differential Equation: dx/dt = 10x-x^2 separable variables so dx/(10x-x^2) = dt

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ 10x-x^2 } = \int\limits dt\]